A +13 nC charge is located at (x, y) = (0 cm, 13 c
Answer #1

for x co ordinate

k ( 4) / 7^2 = k ( 8) / x^2

x = - 9.899 cm

======

for y co ordinate

k (13) / 13^2 = k ( 8) / y^2

y= 10.198 cm

====

comment before rate in case any doubt, will reply for sure.. goodluck